# Equal contribution    Corresponding author

### v3

@@ 4.1 Datasets
To evaluate whether our AGFSync can enhance the performance of text-to-image models across a wide range of prompts, we considered the following benchmarks:
@@
    TIFA [9]: Based on the correct answers to a series of predefined questions. TIFA employs visual question answering (VQA) models to determine if the content of generated images accurately reflects the details of the input text. The benchmark itself is comprehensive, encompassing 4,000 different text prompts and 25,000 questions across 12 distinct categories.
@@
    HPS v2 [28]: Human Preference Score v2 (HPS v2) is an extensive dataset specifically collected to document human preference choices across a variety of images. HPD v2 aims at precisely capturing human preferences for images generated from textual descriptions. HPS v2 provides a benchmark for text-to-image generation models, with a publicly available test dataset comprising 3,200 distinct image captions, covering five types of image descriptions: anime, photo, drawbench, concept-art, and paintings.
@@ 4.2 Hyperparameters
For each given text prompt $\bm{c}$, we let the diffusion model generate $N=8$ samples as backup images for further preference dataset constitution. The weight term in the calculation of CLIP score $\gamma$ is set to $100$, which leads the CLIP Score range between 0 and 100. We also rescale the VQA score and aesthetic score to $0-100$ by multiplying $100$ to original score.
@@
$S(\mathbf{x})=w_{\text{VQA}}s_{\text{VQA}}(\mathbf{x})+w_{\text{CLIP}}s_{\text{CLIP}}(\mathbf{x})+w_{\text{Aesthetic}}s_{\text{Aesthetic}}(\mathbf{x})$ (9)
During DPO alignment stage, we finetune the original diffusion model. For SD v1.4 and SD v1.5 model, the learning rate is 5e-7, the batch size is 128, the output image size is 512*512. For SDXL model, the learning rate is 1e-6, the batch size is 64, the output image size is 1024*1024. We fine-tune the diffusion model for 1000 steps. The random seed is 200 in Fig. 6, Fig. 5, and Fig. 1.
@@ 4.3 Baseline Models and Utilized Models
We evaluated our framework, AGFSync, using Stable Diffusion v1.4 (SD v1.4), Stable Diffusion v1.5 (SD v1.5) [22], and SDXL (SDXL-base-1.0) [17], widely acknowledged in related research as the current leading open-source text-to-image (T2I) models. For prompt construction, we employ ChatGPT (GPT-3.5)[14]. For generating question-answer pairs, we use Gemini Pro[16]. Both are accessed through their official API. In addition, we adopt Salesforce/blip2-flan-t5-xxl for VQA scoring model [12], openai/clip-vit-base-patch16 for evaluating CLIP score [8], Vila [10] for calculating aesthetic score [10], which are consistent with the baseline methods’ settings. We also employ GPT-4 Vision (GPT-4V) [1] to simulate human preference when evaluating the image quality.
@@ 5 Experimental Results
We present the evaluation results of baseline models fine-tuned with our framework AGFSync on benchmarks.
@@ 5.1 Benchmarking Results on HSP v2
5.1.1 Evaluate by CLIP Score and Aesthetic Score
@@
Figure 3: Comparison of the win rates of SD v1.4, SD v1.5 and SDXL with or without our AGFSync on HPS v2. CLIP score (left) and aesthetic score (right).
As in Fig. 3, we test the win rates of models fine-tuned with our method AGFSync and the original models on CLIP score and aesthetic score in the HSP v2 benchmark. The experimental results demonstrate with AGFSync, models have a consistent higher win rate against the original the baseline SD v1.4, SD v1.5, and SDXL models without fine-tuning across a wide range of image descriptions. Notably, following AGFSync fine-tuning, the SDXL model achieves a win rate of 60.5% in CLIP score compared to the original model in the anime category images, and the SD v1.5 model achieves a win rate of 77.4% in aesthetic score for the photo category images. The average win rate of CLIP Score and aesthetic score for the three models increased to 57.2% and 61.6% compared to base ones.
5.1.2 Evaluate by GPT-4 Vision to Simulate Human Preference
(a) Win Rate of SD v1.4
(b) Win Rate of SD v1.5
(c) Win Rate of SDXL
Figure 4: Win rate results of using GPT-4V to evaluate our AGFSync fine-tuned models of SD v1.4, SD v1.5, and SDXL, compared to the original models for general preference (Q1), prompt alignment (Q2), and visual appeal (Q3) on the HSP v2 dataset.
@@ 5.1.2 Evaluate by GPT-4 Vision to Simulate Human Preference
In this study, we explore the efficacy of AGFSync in enhancing image generation models, leveraging the capabilities of GPT-4 Vision (GPT-4V) as reported by OpenAI in 2023 [1] to simulate human preference. Our methodology involves a comparative analysis between images produced by diffusion models before and after the application of AGFSync. These images, accompanied by their respective descriptions, were submitted to GPT-4V for evaluation based on three critical aspects: General Preference (Q1): "Which image do you prefer?"; Prompt Alignment (Q2): "Which image better fits the text description?"; Visual Appeal (Q3): "Disregarding the prompt, which image is more visually appealing?". The evaluation process involved collecting and analyzing the frequency with which images produced by both the original and the trained model are favored under each question category. The results of this comparative analysis are visually represented in Fig. 4, which sequentially displays the performance metrics. The performance reveals that adding AGFSync yields substantial enhancements across all models concerning Q1, Q2 and Q3. Notably, with our AGFSync applied, we achieve an average of 62%, 67%, and 69% win rate of three aspects for the SD v1.4, SD v1.5, and SDXL models respectively. These results demonstrate the effectiveness of AGFSync in enhancing performance under various prompts.
@@ 5.2 Benchmarking Results on TIFA
Table 1: Results of different alignment methods on VQA score and aesthetic score on the TIFA benchmark. Red indicates improvement, while Green indicates a decrease. The best scores for each model type are highlighted in Bold. Column "Sum" denotes the sum of improvements on $s_{\text{VQA}}$ and $s_{\text{Aes.}}$
@@
SDXL No alignment 82.0 60.9 -
@@
  • Note: Due to the code of DreamSync not being open-sourced, the results of SD v1.4 are extracted from the DreamSync paper. We reproduce the results of SD v1.5 and SDXL. We do not apply the Refiner for SDXL. We ensure that the comparisons are fair.
In Tab. 1, We further test our method on TIFA benchmark, highlighting AGFSync’s SOTA performance on VQA score and aesthetic score over other latest SOTA alignment methods. Specifically, we compare three types of alignment methods: training-free approaches capable of modifying outputs without retraining the model, such as StructureDiffusion [7] and SynGen [21]; reinforcement learning (RL)-based methods aimed at improving model outputs, such as DPOK [6] and DDPO [3]; and methods like DreamSync [26], which employ self-training strategy but focus on SFT stage. Given that these baseline methods are all conducted upon SD v1.4, we ensure a fair comparison by using the same version of the SD model as the foundation and employing the same VQA model (BLIP-2) for evaluation. Results reveal that our method AGFSync can simultaneously improve the text fidelity and visual quality of SD v1.4, SD v1.5, and SDXL models. For SD v1.4, AGFSync achieves an improvement of 1.3% of VQA score and 3.3% of aesthetic score, with a sum of 4.6% improvement on the TIFA benchmark, higher than all baseline models. Notice that although DPOK shows a 1.9% improvement on aesthetic score, it reduces the model’s text faithfulness through VQA score. For SD v1.5 and SDXL, our method AGFSync leads to improvements of 1.6% and 1.1% for SD v1.5, 1.3% and 4.3% for SDXL in terms of VQA score and aesthetic score respectively.
@@ 5.3 Ablation Experiment of Multi-Aspect Scoring
Table 2: Results of the ablation study of applied scoring measures. Experiments are conducted with SD v1.5 on TIFA.
@@
Table 3: The number of prompts of different categories and the total number of prompts in our dataset.
Category Count
Natural Landscapes 5733
Cities and Architecture 6291
People 5035
Animals 3089
Plants 4276
Food and Beverages 3010
Sports and Fitness 2994
Art and Culture 2432
Technology and Industry 3224
Everyday Objects 2725
Transportation 4450
Abstract and Conceptual 2575
Total 45834
As depicted in Tab. 3, to validate the efficacy of the three scores we employed for image quality assessment, we conduct a thorough ablation study. We train the SD v1.5 model on preference datasets constructed with different combinations of the three scores, along with PickScore [11]. As in AGFSync, with training model on preference datasets built using a combination of CLIP score, VQA score, and aesthetic score result in the greatest improvement across all three metrics. While other combinations often show a decrease in certain metric rather than a consistent improvement on all metrics.
@@ 5.4 Qualitative Comparison of Faithfulness and Coherence
Figure 5: Text faithfulness comparisons between SDXL and AGFSync (Ours)+SDXL. The top row shows images generated by SDXL, while the second row displays images generated by Ours+SDXL. In the descriptions at the bottom, the red-highlighted portions indicate discrepancies between images generated by SDXL and input prompts.
Figure 6: Comparison of adherence to real-world rules between SDXL and AGFSync (Ours)+SDXL. The first row shows the images generated by SDXL, while the second row displays those generated by Ours+SDXL. The third row illustrates the comparison of details between the two, demonstrating AGFSync’s finer coherence and details.
To visually demonstrate the improvement of AGFSync, we present comparisons of the SDXL model along with AGFSync for the same image description in Fig. 5 and Fig. 6. From Fig. 5, we observe that while the original SDXL model can generate vivid images, there are instances where discrepancies arise between the generated image and the input description. After fine-tuning with AGFSync, the SDXL model generates images that are more consistent with the input description. Additionally, in Fig. 6, we notice that the original SDXL model sometimes generates images that deviate from real-world patterns or physical logic in certain details, whereas the fine-tuned SDXL model with AGFSync performs better in such cases. For instance, in the second column, the girl’s chest generated by SDXL exhibits unrealistic wrinkles, while in the fourth column, the top cake appears suspended during stacking.
5.5 Prompt Utilization Rate Analysis
Figure 7: In our generated dataset constructed with images generated by SDXL, the proportion of prompts that can be filtered out based on varying thresholds.
To compare with DreamSync [26], in Fig. 7, we present the number of prompts that can be selected from our generated dataset as the thresholds for VQA score and aesthetic Score vary. Specifically, these are prompts for which at least one generated image scores above the respective thresholds. With DreamSync’s 0.9 and 0.6 thresholds for two metrics, only 48.8% of prompts satisfying both, meaning a low data conversion efficiency of 48.8% on our dataset. However, AGFSync’s approach merely requires selecting the best and worst images without imposing any threshold constraints, thereby achieving a data conversion efficiency of 100%.
6 Limitations
Firstly, AGFSync relies on existing large language models (LLMs) and aesthetic scoring models, whose performance and accuracy could be influenced by the biases and limitations of the LLMs. Secondly, while we introduce random noise to increase image diversity, this method might lead to a reduction in consistency between some images and their text prompts. In addition, due to the high cost of time or money, we have not adopted LLaVA [13], GPT-4V and latest advanced multimodal large models [18, 13, 4, 34, 29, 33] to generate prompts or QA pairs. In terms of image evaluation, we employ VQA scores, CLIP scores, and aesthetic scores, which may not capture all aspects of image quality.
@@ 7 Conclusions
Our proposed text-to-image generation framework AGFSync, by leveraging Direct Preference Optimization (DPO) and multi-aspect AI feedback, significantly enhances the prompt following ability and image quality regarding style, coherence, and aesthetics. Extensive experiments on the HPSv2 and TIFA benchmark demonstrate that AGFSync outperforms baseline models in terms of VQA scores, CLIP score, aesthetic evaluation. Based on an AI-driven feedback loop, AGFSync eliminates the need for costly human-annotated data and manual intervention, paving the way for scalable alignment techniques.
Appendix
Appendix 0.A Notations
The notations and their descriptions in the paper are shown in Tab. 4.
Table 4: Notations symbols and their descriptions

### v6

@@ 4.1 Datasets
To evaluate whether our AGFSync can enhance the performance of text-to-image models across a wide range of prompts, we consider the following benchmarks:
@@
    TIFA (Hu et al. 2023): Based on the correct answers to a series of predefined questions. TIFA employs visual question answering (VQA) models to determine whether the content of generated images accurately reflects the details of the input text. The benchmark itself is comprehensive, encompassing 4,000 different text prompts and 25,000 questions across 12 distinct categories.
@@
    HPS v2 (Wu et al. 2023): Human Preference Score v2 (HPS v2) is a benchmark designed to evaluate models’ capabilities across a variety of image types. It comprises 3,200 distinct image captions and covers five categories of image descriptions: anime, photo, drawbench, concept-art, and paintings.
@@ 4.2 Hyperparameters
For each given text prompt $\bm{c}$, we let the diffusion model generate $N=8$ samples as backup images for preference dataset construction. In this process, we add Gaussian noise $\bm{n}\sim\mathcal{N}(0,\sigma^{2}\bm{I})$ to the text embedding, where $\sigma$ is set to $0.1$. In the calculation of CLIP score, $\gamma$ is set to $100$, which leads to the CLIP Score range between 0 and 100. We also rescale the VQA score and aesthetic score to $0-100$ by multiplying the original score by $100$.
@@
$S(x)=0.35s_{\text{VQA}}(x)+0.55s_{\text{CLIP}}(x)+0.1s_{\text{Aes.}}(x)$ (9)
During the DPO alignment stage, we finetune the original diffusion model. For the SD v1.4 and SD v1.5 models, the learning rate is 5e-7, the batch size is 128, the output image size is $512\times 512$. For the SDXL-base model, the learning rate is 1e-6, the batch size is 64, the output image size is $1024\times 1024$. We finetune the diffusion model for 1,000 steps. The random seed is set to 200 in Fig. 4(b) and Fig. 4(a).
@@ 4.3 Baseline Models and Utilized Models
We evaluate AGFSync using Stable Diffusion v1.4 (SD v1.4), Stable Diffusion v1.5 (SD v1.5) (Rombach et al. 2022b), and Stable Diffusion XL Base 1.0 (SDXL-base) (Podell et al. 2023), widely acknowledged in related research as the current leading open-source text-to-image (T2I) models. For prompt construction, we employ ChatGPT (GPT-3.5) (OpenAI 2023). For generating Q&A pairs, we use Gemini Pro (Pichai and Hassabis 2023). Both are accessed through their official API. In addition, we adopt Salesforce/blip2-flan-t5-xxl for VQA scoring model (Li et al. 2023), openai/clip-vit-base-patch16 for evaluating CLIP score (Hessel et al. 2021), Vila (Ke et al. 2023) for calculating aesthetic score (Ke et al. 2023), which are consistent with the baseline methods’ settings. We also employ GPT-4 Vision (GPT-4V) (Achiam et al. 2023) to simulate human preferences when evaluating the image quality.
@@ 5.1 Benchmarking Results on HSP v2
Evaluate by CLIP Score and Aesthetic Score
As in Fig. 2, we test the win rates of models finetuned with our method AGFSync against the original models on the CLIP score and aesthetic score in the HSP v2 benchmark. The experimental results show that with AGFSync, models consistently achieve win rates exceeding 50% across all image categories and both evaluation metrics, CLIP score and aesthetic score, compared to the original baseline SD v1.4, SD v1.5, and SDXL-base models without finetuning. Notably, after AGFSync finetuning, the SDXL-base model not only achieves a win rate of 60.5% in the CLIP score compared to the original model in the anime category images, but also achieves a win rate of 77.4% in the aesthetic score for the same category images. The average win rate of the CLIP score and aesthetic score for the three models increases to 57.2% and 61.6%, respectively, compared to base ones.
@@
Figure 2: Comparison of the win rates of SD v1.4, SD v1.5 and SDXL-base with or without our AGFSync on HPS v2. CLIP score (left) and aesthetic score (right).
Evaluate by GPT-4 Vision to Simulate Human Preference
@@
In this study, we explore the efficacy of AGFSync in enhancing image generation models, leveraging the capabilities of GPT-4 Vision (GPT-4V) as reported by OpenAI in 2023 (Achiam et al. 2023) to simulate human preferences. Our methodology involves: (1) a comparative analysis of images generated by various diffusion models before and after the application of AGFSync; and (2) a comparative analysis of images generated by the SD v1.4 model after applying AGFSync or other alignment methods. These images, accompanied by their respective descriptions, are submitted to GPT-4V for evaluation based on three critical aspects: General Preference (Q1): “Which image do you prefer?”; Prompt Alignment (Q2): “Which image better fits the text description?”; Visual Appeal (Q3): “Disregarding the prompt, which image is more visually appealing?”.
Table 1: Win rate results of using GPT-4V to evaluate our finetuned models based on SD v1.4, SD v1.5, and SDXL-base, compared to the original models and models aligned with DDPO, Structured Diffusion, and SynGen (only on SD v1.4), for general preference (Q1), prompt alignment (Q2), and visual appeal (Q3) on the HSP v2 dataset.
Test Model Method General Faithful Aesthetic
SD v1.4 vs Original 62% 58% 65%
vs DDPO 68% 78% 82%
vs Structured Diffusion 64% 70% 79%
vs SynGen 61% 58% 58%
SD v1.5 vs Original 68% 67% 65%
SDXL-base vs Original 62% 69% 76%
The evaluation process involves collecting and analyzing the frequency with which images produced by both the original and the finetuned model are favored under each question category. The results of the comparative analysis of images generated by various diffusion models before and after the application of AGFSync are presented in Table 1, which sequentially displays the performance metrics. The performance reveals that adding AGFSync yields substantial enhancements across all models concerning Q1, Q2, and Q3. Notably, with our AGFSync applied, we achieve an average of 62%, 67%, and 69% win rates across the three aspects for the SD v1.4, SD v1.5, and SDXL-base models respectively. The results of the comparative analysis of images generated by the SD v1.4 model after applying AGFSync or other alignment methods are also presented in Table 1. The performance shows AGFSync consistently outperforms other baselines (DDPO, Structured Diffusion, and SynGen) applied to the SD v1.4 in all dimensions, achieving average win rates of 64.2%, 68.4%, and 72.8% respectively, when evaluated by GPT-4V on images generated from the HPS v2. These results demonstrate the effectiveness of AGFSync in enhancing performance under various prompts.
Human Evaluation of GPT-4V Judgments
To validate the efficacy of GPT-4V for image evaluation and address potential biases in AI assessments, we compare its consistency with human evaluations. We randomly select 9 pairs of images generated by the AGFSync that are favored by GPT-4V. A total of 58 graduate students from China participate in the evaluation. Each participant assesses each image pair based on the criteria Q1, Q2, Q3. Each image pair is independently rated on these three dimensions, resulting in 27 questions per participant (9 image pairs $\times$ 3 dimensions). For Q1, there is 78% agreement between GPT-4V and human evaluations, for Q2, 83%, and for Q3, 70%. All dimensions show agreement rates above 50%, indicating that GPT-4V’s evaluations align closely with human preferences and confirming its reliability as a tool for reducing individual biases and maintaining objectivity in image evaluation.
@@ 5.2 Benchmarking Results on TIFA
In Table 2, we further test our method on the TIFA benchmark, highlighting AGFSync’s SOTA performance on VQA score and aesthetic score over other latest SOTA alignment methods. Specifically, we compare three types of alignment methods: training-free approaches capable of modifying outputs without retraining the model, such as StructureDiffusion (Feng et al. 2023) and SynGen (Rassin et al. 2023); reinforcement learning (RL)-based methods aimed at improving model outputs, such as DPOK (Fan et al. 2024) and DDPO (Black et al. 2023); and methods like DreamSync (Sun et al. 2023), which employ self-training strategy but focus on SFT stage. Given that these baseline methods are all based on SD v1.4, we ensure a fair comparison by using the same version of the SD model as the foundation and employing the same VQA model (BLIP-2) for evaluation. Results reveal that our method AGFSync can simultaneously improve the text fidelity and visual quality of SD v1.4, SD v1.5, and SDXL-base models. For SD v1.4, AGFSync achieves an improvement of 1.3% of VQA score and 3.3% of aesthetic score, with a total improvement of 4.6% on the TIFA benchmark, higher than all baseline models. Note that although DPOK shows a 1.9% improvement on aesthetic score, it reduces the model’s text faithfulness through VQA score. For SD v1.5 and SDXL-base, our method AGFSync leads to improvements of 1.6% and 1.1% for SD v1.5, 1.3% and 4.3% for SDXL-base on VQA score and aesthetic score respectively, which are both higher than the results achieved by DreamSync finetuned using self-training SFT.
Table 2: Results of different alignment methods on VQA score and aesthetic score on the TIFA benchmark. Red indicates improvement, while Green indicates a decrease. The best scores for each model type are in Bold. Column “Sum” denotes the sum of improvements on $s_{\text{VQA}}$ and $s_{\text{Aes.}}$
@@
SDXL-base No alignment 82.0 60.9 -
DreamSync 83.1 ($+$1.1) 64.1 ($+$3.2) $+$4.3
@@
5.3 Experiment of Comparing the Dataset Quality between MJHQ-30K and AGFSync
MJHQ-30K is a benchmark dataset used for automatically evaluating the aesthetic quality of models (Li et al. 2024). It consists of high-quality images curated from Midjourney, covering 10 common categories, with each category containing 3,000 samples. MJHQ-30K can also serve as a training dataset for general SFT. To compare the quality of the preference dataset built using AGFSync with MJHQ-30K, we finetune SD v1.4, SD v1.5 and SDXL-base using MJHQ-30K and compare their performance against the SD v1.4, SD v1.5 and SDXL-base finetuned with AGFSync. As shown in Table 3, AGFSync applied to SD v1.4, SD v1.5 and SDXL-base achieve superior improvements in text alignment. Although finetuning SD v1.4 and SD v1.5 with the MJHQ-30K dataset results in the highest improvement in aesthetic scores, this is because the images in MJHQ-30K is generated by Midjourney, which have much higher aesthetic quality than those generated by SD v1.4 and SD v1.5 for self-training. When finetuning SDXL-base with MJHQ-30K, the improvement in aesthetic scores is less pronounced compared to AGFSync, demonstrating the effectiveness of the preference dataset constructed using AGFSync.
Table 3: SD v1.4, SD v1.5 and SDXL-base’s results of general SFT setting on MJHQ-30K compared to AGFSync.
Model Alignment $s_{\text{VQA}}$ $s_{\text{Aes.}}$ Sum
SD v1.4 No alignment 76.6 44.6 -
MJHQ-30K+SFT 77.6 (+1.0) 48.3 (+3.7) +4.7
AGFSync (Ours) 77.9 (+1.3) 47.9 (+3.3) +4.6
SD v1.5 No alignment 77.1 48.0 -
MJHQ-30K+SFT 78.3 (+1.2) 49.3 (+1.3) +2.5
AGFSync (Ours) 78.7 (+1.6) 49.1 (+1.1) +2.7
SDXL-base No alignment 82.0 60.9 -
MJHQ-30K+SFT 82.6 (+0.6) 61.1 (+0.2) +0.8
AGFSync (Ours) 83.3 (+1.3) 65.2 (+4.3) +5.6
5.4 Experiment of Gaussian Noise for Diversity
To demonstrate that adding Gaussian noise $n$ to a given condition $c$ during the generation of $N$ candidate images significantly enhances the diversity of the candidates, we conduct an experiment as shown in the Fig. 3. We utilize a consistent text input “wild animal” to generate four images and systematically adjust the weight of the noise $n$. By comparing the generated images under different noise weights, we observe significant changes in the variety of animal species with the increasing weight of the noise (Fig. 3). Notably, when the noise weight is set to 1, the images exhibit the most diverse range of animal species. This finding supports our hypothesis that the introduction of Gaussian noise effectively expands the coverage of the conditional input, thus increasing the exploration space of the model and significantly enhancing the diversity of the generated images.
Figure 3: Impact of noise on image diversity. With the numbers on the left side of the images indicating the increasing weight of noise, four images were generated using the same text input “wild animal”.
@@ 5.5 Ablation Experiment of Multi-Aspect Scoring
As depicted in Table 4, to validate the efficacy of the three scores that we employ for image quality assessment, we conduct a thorough ablation study. We train the SD v1.5 model on preference datasets constructed with different combinations of the three scores, along with PickScore (Kirstain et al. 2024). As in AGFSync, with training model on preference datasets built using a combination of CLIP score, VQA score, and aesthetic score result in the greatest improvement across all three metrics. While other combinations often show a decrease in certain metrics rather than a consistent improvement on all metrics.
Table 4: Results of applied scoring measures. Experiments are conducted with SD v1.5 on TIFA.
@@ 5.6 Qualitative Comparison of Faithfulness and Coherence
Figs. 4(a) and 4(b) compare images generated by SDXL-base and AGFSync using identical prompts. While SDXL-base generates vivid images, they sometimes deviate from input descriptions (Fig. 4(a)) or contain unrealistic details (Fig. 4(b)). For example, SDXL-base produces unnatural wrinkles on a girl’s chest and physically impossible floating cakes. After finetuning with AGFSync, the generated images show improved consistency with prompts and better adherence to real-world physics.
(a) Text faithfulness comparison
(b) Adherence to real-world rules
Figure 4: Comparison between SDXL-base and AGFSync (Ours)+SDXL-base. (a) Red-highlighted text indicates discrepancies with input prompts. (b) Third row compares details, showing AGFSync’s improved coherence and detail.
@@ 6 Conclusions
This paper introduces a text-to-image generation framework AGFSync. By leveraging Direct Preference Optimization (DPO) and multi-aspect AI feedback, AGFSync significantly enhances the prompt following ability and image quality regarding style, coherence, and aesthetics. Extensive experiments on the HPSv2 and TIFA benchmark demonstrate that AGFSync outperforms baseline models in terms of VQA scores, CLIP score, aesthetic evaluation. Based on an AI-driven feedback loop, AGFSync eliminates the need for costly human-annotated data and manual intervention, paving the way for scalable alignment techniques.
Acknowledgments
This work was supported by the National Key R&D Program of China (No. 2023YFB3309000), the National Natural Science Foundation of China under Grants U2241217, 62473027 and 62473029.
Appendix A Notations
The notations and their descriptions in the paper are shown in Table 5.
Table 5: Notations symbols and their descriptions.